OpenStudy (anonymous):

3 1/2 - 2/3 need to see the work

OpenStudy (anonymous):

\[3 + 1/2 = 7/2 \] so to find the common denominator of 7/2 and 2/3 in order to subtract you can multiply each fraction by one in a clever way to have both the denominators equal. \[7/2 * 3/3\] and \[2/3 * 2/2\] now you have your equation where 21/6 = 3 1/2 and 2/3 = 4/6 therefore you may subtract as follows\[(21-4)/6\]

OpenStudy (anonymous):

isnt tht 2 5/6?

OpenStudy (anonymous):

yes 17/6 = 2 and 5/6 if all the above turns out correct that would be your final answer in compound fraction form.

OpenStudy (anonymous):

but on this practice sheet it says the awnser is 1 1/6

OpenStudy (anonymous):

here is the question

OpenStudy (anonymous):

erica bought 3 1/2 yards of fabric. if she uses 2/3 of the fabric to make a curtain, how much will she have left?

OpenStudy (anonymous):

oh okay. she is using 2/3 of the 3.5 yards not 2/3 yards

OpenStudy (anonymous):

so here you can take your 3 and 1/2 or 7/2 yards and divide it by three as follows \[7/2/3 = 7/2*3 = 7/6 \] that will give you one third and we're looking for 1/3 of the quantity. because 2/3 was used.

OpenStudy (anonymous):

7/6 = 1 and 1/6

OpenStudy (anonymous):

divide by three because you are looking for one third of the fabric. 2/3 of it was used so your full amount (1) minus 2/3 is 1/3.